I have two sister. I am tree times older than one of my sisters and six years older than the other sister.If one of my sisters is half the age of the other sister,how old am I?
Answer by ankor@dixie-net.com(22740)

Let a = The speaker's age.
let b = one sister's age
let c = the other sister's age
:
Write an equation for each statement
:
I have two sisters.
I am three times older than one of my sisters,
a = 3b
and six years older than the other sister.
a = c+6
If one of my sisters is half the age of the other sister,
b = .5c
how old am I?
;
From the first two equations; a = a, therefore
3b = c+6
replace b with .5c
3(.5c) = c + 6
1.5c = c + 6
1.5c - c = 6
.5c = 6
c = 12
then
a = 12 + 6
a = 18 yrs old
:
:
Check this by finding the age of the other sister
b = .5(12)
b = 6
and
a = 3(6) = 18